Find y(t) solution of the initial value problem 3t y? Y – 6 y3 – 4t² = 0, y(1) = 1, g(1)=1, t > 0.
Answer #1

S y의 3y -62-H= - 69 나 3y 39 - 35 바 2년 Comparing with y=오 - 2 른 3멍드 Integrating factor e Sadh S른다 2 e logt logt t2 e 른 고 가브 roGerend suluhion is, 닛 1 2 놂 (10919 3y 그 (lugt) T 에 39 누르 3y 35 난 Clog t 3가282.